if (x+y)^2=40 and (x-y)^2=25, evaluate 2x^2-40xy+2y^2
Answer by Alan3354(69443)

if (x+y)^2=40 and (x-y)^2=25, evaluate 2x^2-40xy+2y^2
------------
x^2 + 2xy + y^2 = 40
x^2 - 2xy + y^2 + 25
---------------------------- Subtract
4xy = 15
40xy = 150
===================
2x^2-40xy+2y^2 = 40 + 25 - 150
--> -85